you drive your car in a straight line at 15 m/s for 10 minutes then at 25 m/s for another 10 minutes. is your average speed for entire trip more than, less than or equal to 20 ms

1
Expert's answer
2021-09-12T19:01:34-0400

By definition, the average speed is given as total distance divided by total time:


"v_{av} = \dfrac{d_1+d_2}{t_1 + t_2}"

where "d_1,d_2" are the parts of total distance covered during times "t_1 = 10min = 600s,t_2 = 10min = 600s" with speeds "v_1 = 15m\/s, v_2 = 25m\/s" respectively.

For the first half of the journey, obtain:


"d_1 = v_1t_1"

For the second half:


"d_2 = v_2t_2"

Finally, obtain:


"v_{av} = \dfrac{v_1t_1+v_2t_2}{t_1 + t_2} = \dfrac{15\cdot 600+25\cdot 600}{600 + 600} = 20m/s"

Answer. equal to 20 m/s.